# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
$11^{\prime*}$ $B^{-}\to T^{0}_{c}\kappa^{-}$ $-\sqrt{\frac{2}{3}}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\overline{\kappa}{}^{0}$ $-\sqrt{2}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$
@@
$14^{\prime}$ $B^{-}\to T^{\prime 0}_{c}\kappa^{-}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}-C^{\prime}_{S}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+3C^{\prime}_{TS,1})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\overline{\kappa}{}^{0}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}+C^{\prime}_{S}-C^{\prime}_{ST,1}-2C^{\prime}_{ST,2})$
@@
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\sigma$ $-\frac{\sqrt{3}s_{\phi}}{2\sqrt{2}}(2C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1}+E^{\prime}_{\bar{B}T})+\frac{\sqrt{3}c_{\phi}}{2}(C^{\prime}_{TS,1}-E^{\prime}_{\bar{B}T})$
@@
$16^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c}a^{-}$ $-\frac{1}{2}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-4E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
@@
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $-\frac{1}{2\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-4E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $-\frac{c_{\phi}}{2\sqrt{2}}(2C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}+3C^{\prime}_{ST,2}+8C^{\prime}_{ST,3}+C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+5E^{\prime}_{\bar{B}T})$
$-\frac{s_{\phi}}{2}(C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})$
@@
$-\frac{c_{\phi}}{2}(C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})$
@@
$2\Gamma(\overline{B}{}^{0}\to T^{0}_{c\bar{s}}\overline{\kappa}{}^{0})$ $=$ $\Gamma(\overline{B}{}^{0}\to T^{-}_{c}a_{0}^{+})=2\left|\frac{V_{ud}}{V_{us}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to T^{0}_{cs}\kappa^{0})$ (71)
@@ VI $B^{-}_{c}\to T\overline{T}$ decays
From the above discussion, we note that $T^{++}_{c\bar{s}}$ and other doubly charged states do not participate in $\overline{B}\to\overline{D}T$, $\overline{B}\to TP$ and $\overline{B}\to TS$ decays, as required by charge conservation. It is certainly welcome to have modes including them besides those in $\overline{B}\to D\overline{T}$ decays, where $T^{++}_{c\bar{s}}$ was discovered. In this section, we consider $B^{-}_{c}$ decays, which are kinematically allowed to produce a $T\overline{T}$ pair. As we shall see $T^{++}_{c\bar{s}}$ and other doubly charged states can be produced in $B^{-}_{c}\to T\overline{T}$ decays.
@@ VI.1 Topological amplitudes in $B^{-}_{c}\to T\overline{T}$ decays
For notational simplicity, we do not distinguish the topological amplitudes in scenarios I and II. We have seven topological amplitudes, namely $C_{T\bar{T}}$, $C_{\bar{T}T,1}$, $A_{T\bar{T}}$, $C_{T}$, $C_{\bar{T}}$, $C_{\bar{T}T,2}$ and $A_{\bar{T}T}$, in $B^{-}_{c}\to T\overline{T}$ decays, where $C$ and $A$ are internal $W$ and annihilation diagrams, respectively. The subscripts in the topological amplitudes denote the tetraquark(s) where the $q^{i}\bar{q}_{j}$, from the $b\to cW^{-}\to cq^{i}\bar{q}_{j}$ transition, contribute to. The topological amplitudes are depicted in Fig. 8. Diagrams (a) to (c) contribute to modes with flavor exotic $T$ and $\overline{T}$.
@@
$(3^{*},\bar{2}^{*})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(6,\bar{5})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $-\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(3^{*},\bar{9})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(2C_{T}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(6,\bar{7})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $-\frac{1}{2}(2C_{T}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{\bar{T}T})$
@@
$(9,\bar{1}^{*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(2C_{\bar{T}}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
@@
$(4^{*},\bar{5})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
$(5,\bar{1}^{*})$ $B^{-}_{c}\to T^{+}_{c}\overline{T}^{\,--}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
@@
$(4^{*},\bar{7})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$

### v3

@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
$11^{\prime*}$ $B^{-}\to T^{0}_{c}\kappa^{-}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\overline{\kappa}{}^{0}$ $\sqrt{2}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$
@@
$14^{\prime}$ $B^{-}\to T^{\prime 0}_{c}\kappa^{-}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}-C^{\prime}_{S}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+3C^{\prime}_{TS,1})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\overline{\kappa}{}^{0}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{T}+C^{\prime}_{S}-C^{\prime}_{ST,1}-2C^{\prime}_{ST,2})$
@@
$14^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime 0}_{c}\sigma$ $\frac{\sqrt{3}s_{\phi}}{2\sqrt{2}}(2C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1}+E^{\prime}_{\bar{B}T})+\frac{\sqrt{3}c_{\phi}}{2}(C^{\prime}_{TS,1}-E^{\prime}_{\bar{B}T})$
@@
$16^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c}a^{-}$ $\frac{1}{2}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-4E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
@@
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $\frac{1}{2\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-4E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $\frac{c_{\phi}}{2\sqrt{2}}(2C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}+3C^{\prime}_{ST,2}+8C^{\prime}_{ST,3}+C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+5E^{\prime}_{\bar{B}T})$
$\frac{s_{\phi}}{2}(C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})$
@@
$\frac{c_{\phi}}{2}(C^{\prime}_{TS,1}+4E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})$
@@
[…]
@@ VI $B^{-}_{c}\to T\overline{T}$ decays
From the above discussion, we note that $T^{++}_{c\bar{s}}$ and other doubly charged states do not participate in $\overline{B}\to\overline{D}T$, $\overline{B}\to TP$ and $\overline{B}\to TS$ decays, as required by charge conservation. It is preferable to have other modes to include them, besides $\overline{B}\to D\overline{T}$ decays, where $T^{++}_{c\bar{s}}$ was discovered. In this section, we consider $B^{-}_{c}$ decays, which are kinematically allowed to produce a $T\overline{T}$ pair. As we shall see $T^{++}_{c\bar{s}}$ and other doubly charged states can be produced in $B^{-}_{c}\to T\overline{T}$ decays.
@@ VI.1 Topological amplitudes in $B^{-}_{c}\to T\overline{T}$ decays
For notational simplicity, we do not distinguish the topological amplitudes in scenarios I and II. We have seven topological amplitudes, namely $C_{T\bar{T}}$, $C_{\bar{T}T,1}$, $A_{T\bar{T}}$, $C_{T}$, $C_{\bar{T}}$, $C_{\bar{T}T,2}$ and $A_{\bar{T}T}$, in $B^{-}_{c}\to T\overline{T}$ decays, where $C$ and $A$ are internal $W$ and annihilation diagrams, respectively. The subscripts in the topological amplitudes denote the open-charmed tetraquark(s), where the $q^{i}\bar{q}_{j}$, from the $b\to cW^{-}\to cq^{i}\bar{q}_{j}$ transition, contribute to. The topological amplitudes are depicted in Fig. 8. Diagrams (a) to (c) contribute to modes with flavor exotic $T$ and $\overline{T}$.
@@
Comparing Fig. 8 to Figs.  6, we see that the topological diagrams in $B^{-}_{c}\to T\overline{T}$ decay have an additional $q\bar{q}$ line produced. According to the OZI rule, one expects some suppression factors. But the diagrams in $B^{-}_{c}\to T\overline{T}$, see Fig. 8, and $\overline{B}\to TP$ decays, see Figs.  6, are rather different, we cannot offer a reliable comparison of these rates on this basis.
@@
$(3^{*},\bar{2}^{*})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(6,\bar{5})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(3^{*},\bar{9})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(2C_{T}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(6,\bar{7})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{2}(2C_{T}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{\bar{T}T})$
@@
$(9,\bar{1}^{*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(2C_{\bar{T}}-C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
@@
$(4^{*},\bar{5})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
$(5,\bar{1}^{*})$ $B^{-}_{c}\to T^{+}_{c}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
@@
$(4^{*},\bar{7})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$